OpenStudy (anonymous):

\[\int\limits_{0}^{1}(3x^2+2)^3 xdx\]

myininaya (myininaya):

let u=3x^2+2

OpenStudy (anonymous):

How do I evaluate that then?

myininaya (myininaya):

du/dx=6x so du=6xdx or du/6=xdx

myininaya (myininaya):

if x=1, then u=3(1)^2+2=3+2=5 upper limit if x=0, then u=3(0)^2+2=0+2=2 lower limit

myininaya (myininaya):

\[\int\limits_{2}^{5}(1/6)*u^{3}du\]